OpenStudy (anonymous):

who wants to rationalize some expressions :D? 3xy/xy+x = 3?

OpenStudy (anonymous):

not = but the answer is 3 after cancling

OpenStudy (anonymous):

oo we're not even rationalizing. I meant just simplify rational expressions.

OpenStudy (anonymous):

which mean x =0 ?

OpenStudy (anonymous):

3 + x = 3 ?

OpenStudy (anonymous):

factor then cancle

OpenStudy (anonymous):

3xy/xy+x xy+x factors to just x(y) right?

OpenStudy (dumbcow):

is the denominator (xy+x)

OpenStudy (anonymous):

which would then cancle with the xy on top leaving only 3. yes dumb

OpenStudy (dumbcow):

\[\frac{3xy}{xy+x} = \frac{3xy}{x(y+1)} = \frac{3y}{y+1}\]

OpenStudy (dumbcow):

only the x cancels

OpenStudy (anonymous):

hummmm..

OpenStudy (anonymous):

Oh you ment 3xy/(xy+x) :)

OpenStudy (dumbcow):

when you factor something out, divide by each term xy/x = y x/x = 1

OpenStudy (anonymous):

dumb would you look at another?

OpenStudy (dumbcow):

sure